OpenStudy (anonymous):

A body was found at 10 a.m. outdoors on a day when the temperature was 40oF. The medical examiner found the temperature of the body to be 80oF. What was the approximate time of death? Use Newton's law of cooling, with k = 0.1947.

OpenStudy (anonymous):

can anyone help me out on this??

OpenStudy (anonymous):

We have: \[ T=S+(T_0-S)e^{-kt} \]Right? Let's solve for \(t\): \[ \frac{T-S}{T_0-S}=e^{-kt} \]So we have: \[ \ln\left(\frac{T-S}{T_0-S}\right)=-kt \]And, finally: \[ -\frac{1}{k}\ln\left(\frac{T-S}{T_0-S}\right)=t \]A body's normal temperature is: \(T_0\approx98.6\). Try using that.

OpenStudy (anonymous):

right i know the formula

OpenStudy (anonymous):

just having trouble with times

OpenStudy (anonymous):

I have 5am, 8am, 9:30, and 9:45 am

OpenStudy (anonymous):

How many hours and minutes have you calculated have passed?

OpenStudy (anonymous):

I'm thinking 8am

OpenStudy (anonymous):

What did you get for \(t\)?

OpenStudy (anonymous):

around 2

OpenStudy (anonymous):

hours

OpenStudy (anonymous):

I would agree. So, then 10-2 is 8AM.

OpenStudy (anonymous):

you agree with? t=2hours?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

Sure thing.